A basket contains 12 ripe and 7 unripe mangoes. If four mangoes are taken from the basket one after the other, how many samples space we have in this experiment?

1
Expert's answer
2022-03-03T18:01:03-0500

Number of samples: "N=C_4^0+C_4^1+C_4^2+C_4^3+C_4^4=1+4+6+4+1=16."